OpenStudy (anonymous):

This is the hardest question yet! During the basketball season, Diane took 134 shots and made about 56% of them. a. How many shots did Diane make? b. The team made a total of 498 shots. What percent of the team’s made shots did Diane make? 24 shots; 15.1% 753 shots; 27% 240 shots; 27% 75 shots; 15.1% Thanks Again every one +D

OpenStudy (anonymous):

\(\large \frac{made \space shots}{134}=\frac{56}{100}\) Can you solve the proportion?

OpenStudy (lukebluefive):

To get the number of shots, multiply the total number of shots by the percentage that she made: \[134 \times 0.56 = ?\] Round the answer down to the nearest integer. Take this number and divide it by 498, the total number of shots her team made, to get her overall percentage of shots made for her team.

OpenStudy (anonymous):

72.36 = 72 72 / 498 = .1445 ???

OpenStudy (anonymous):

@LukeBlueFive

OpenStudy (anonymous):

Multiply a decimal fraction by 100 to get percent.

OpenStudy (lukebluefive):

Yes, 0.1445 = 14.45%

OpenStudy (lukebluefive):

But 134 x 0.56 ≠ 72.36. Maybe double-check your math?